OpenStudy (anonymous):

Could anyone help me come up with a game?? in this assignment, you will be designing your own game. The game should be the type of game that you would play at a carnival, amusement park or casino . It cannot be a game that already exists — you must create a unique game and you must be able to explain the probability of your game

OpenStudy (anonymous):

a dice game?

OpenStudy (anonymous):

Its got to be a game that doesn't exist though...

OpenStudy (anonymous):

you can make up a dice game i bet make it simple so the probabilities will not be too complicated

OpenStudy (anonymous):

unless you want it to be complicated...

OpenStudy (anonymous):

Not trying to make it complicated, i already suck at math as it is.. :(

OpenStudy (anonymous):

that is why i suggested using dice, those probabilities are easy to get

OpenStudy (anonymous):

Yea true..

OpenStudy (anonymous):

do you know those? like the probability you get 7 on two dice? or roll doubles?

OpenStudy (anonymous):

I struggle with it. :(

OpenStudy (anonymous):

they are not too bad to understand there are 6 ways to roll doubles, double 1, double 2, double 3, double 4, double 5, double 6 and \(6\times 6=36\) possible rolls total, so the probability you get doubles is \(\frac{6}{36}=\frac{1}{6}\)

OpenStudy (anonymous):

clear enough?

OpenStudy (anonymous):

Yeah so.. what could i make my game? like the name and explaining how to play it?

OpenStudy (anonymous):

you can name it anything you like as for making it up, you should make it simple, something like " the gambler pays $10 to roll the dice and wins the amount shown on the dice if he rolls doubles, loses his $10 if she does not" or something like that, something not complicated

OpenStudy (anonymous):

Ok thanks so much, i got a good idea now! :)

OpenStudy (anonymous):

I would fan you but i already have :)

OpenStudy (anonymous):

ok good, make up your own and if you get stuck with the probabilities come back and ask

OpenStudy (anonymous):

ok thanks so muchh!!

OpenStudy (anonymous):

thanks! i don't need you to fan me twice, i am cool enough yw